OpenStudy (anonymous):

Using Gauss' Law, find the electric potential a distance s above a infinite line of charge /lambda

OpenStudy (anonymous):

i have \[V = -\int\limits_{\inf}^{s}E*dl = -C \int\limits_{\inf}^{s} {ds}/{s}\] where C is the constants, but this gives me potential of ln(inf) - ln(s) = inf, where did i go wrong?